f(X) = x^2 + 2. How is this related to f(x) = -x^2 - 1? I think that it is reflected up 3, but my friends say it is reflected down 3. Can you help? Thank you.
Answer by richard1234(7193)

The graph of f(x) = x^2 + 2 can be obtained by reflecting the graph of f(x) = -x^2 - 1 across the line y = 1/2.